# Seeking to Collide: Online Safety-Critical Scenario Generation for Autonomous Driving with Retrieval Augmented Large Language Models

### v1

@@ Abstract
Simulation-based testing is crucial for validating autonomous vehicles (AVs), yet existing scenario generation methods either overfit to common driving patterns or operate in an offline, non-interactive manner that fails to expose rare, safety-critical corner cases. In this paper, we introduce an online, retrieval-augmented large language models (LLMs) framework for generating safety-critical driving scenarios. Our method first employs an LLM-based behavior analyzer to infer the most dangerous intent of the background vehicle from the observed state, then queries additional LLM agents to synthesize feasible adversarial trajectories. To mitigate catastrophic forgetting and accelerate adaptation, we augment the framework with a dynamic memorization and retrieval bank of intent–planner pairs, automatically expanding its behavioral library when novel intents arise. Evaluations using the Waymo Open Motion Dataset demonstrate that our model reduces the mean minimum time-to-collision from 1.62 to 1.08 s and incurs a 75% collision rate, substantially outperforming baselines.
@@ III-C LLM-based Analyzer for Safety-critical Behavioral Intent
Analysis Requirements includes the behavior library, rules for automatically updating the library and the five required analysis steps. Drawing inspiration from the chain-of-thought prompting method [25], the LLM is guided through the five-step reasoning process. It first interprets the current states of both vehicles, then contextualizes these within the map topology. Based on the ego vehicle’s inferred intent, the model evaluate which behavior from the library would most likely result in a high-risk interaction. Finally, it assigns an appropriate acceleration $y_{acc}$ to simulate this behavior in downstream scenario generation.
@@ III-D3 Simulation-Guided Code Refinement
The generated trajectory $\widehat{\mathbf{Y}}^{bac}$ is further evaluated in the simulation testing of the ego vehicle. If the synthesized trajectory fails to induce meaningful safety-critical events, it is sent back to the Code Modifier, a refinement module that adjusts the original code block $f_{\phi}$ based on the feedback. This refinement continues until the synthesized trajectory meets the predefined risk threshold, ensuring the resulting scenario is genuinely adversarial and behaviorally consistent.
@@ III-E1 Intent‐based Retrieval
At time $t$, given a newly inferred intent
$\phi_{t}^{\prime}=g_{\theta}(S_{1:t}^{ego^{\prime}},\,\mathbf{S}_{1:t}^{bac^{\prime}})\,,$
we search $\mathcal{M}$ for the closest stored label:
@@
where $d(\cdot,\cdot)$ is a suitable label‐similarity metric, e.g., the cosine similarity between the text embeddings of intent labels. If $d\bigl{(}\phi_{t}^{\prime},\phi_{j^{*}}\bigr{)}\;\leq\;\epsilon_{\mathrm{ret}}$, we then retrieve and reuse the existing planner:
@@ V Conclusion
There are several directions for further enhancing interactive scenario generation. First, the uncertainty within the intent inference can be integrated. Second, extending the memory bank to support hierarchical intents and multi-agent coordination may yield more complex adversarial interactions.

### v2

@@ Abstract
Simulation-based testing is crucial for validating autonomous vehicles (AVs), yet existing scenario generation methods either overfit to common driving patterns or operate in an offline, non-interactive manner that fails to expose rare, safety-critical corner cases. In this paper, we introduce an online, retrieval-augmented Large Language Models (LLMs) framework for generating safety-critical driving scenarios. Our method first employs an LLM-based behavior analyzer to infer the most dangerous intent of the background vehicle from the observed state, then queries additional LLM agents to synthesize feasible adversarial trajectories. To mitigate catastrophic forgetting and accelerate adaptation, we augment the framework with a dynamic memorization and retrieval bank of intent-planner pairs, automatically expanding its behavioral library when novel intents arise. Evaluations using the Waymo Open Motion Dataset demonstrate that our model reduces the mean minimum time-to-collision from 1.62 to 1.08 s and incurs a 75% collision rate, substantially outperforming baselines.
@@ III-C LLM-based Analyzer for Safety-critical Behavioral Intent
Analysis Requirements includes the behavior library and the five required analysis steps. The behavior library is extensible and can be updated by LLM analyzer automatically, which is described in Section III-E. Drawing inspiration from the chain-of-thought prompting method [25], the LLM is guided through the five-step reasoning process. It first interprets the current states of both vehicles, then contextualizes these within the map topology. Based on the ego vehicle’s inferred intent, the model evaluate which behavior from the library would most likely result in a high-risk interaction. Finally, it assigns an appropriate acceleration $y_{acc}$ to simulate this behavior in downstream scenario generation.
@@ III-D3 Simulation-Guided Code Refinement
The generated trajectory $\widehat{\mathbf{Y}}^{bac}$ is further evaluated in the simulation testing of the ego vehicle. If the synthesized trajectory fails to induce meaningful safety-critical events, it is sent back to the Code Modifier, a refinement module that adjusts the original code block $f_{\phi}$ based on the feedback. This refinement continues until the synthesized trajectory meets the predefined risk threshold, ensuring the resulting scenario is genuinely adversarial and behaviorally consistent. The LLM-based code generator and the modifier are both implemented by prompting a general-purpose LLM with carefully engineered instructions.
@@ III-E1 Intent‐based Retrieval
At time $t$, given a newly inferred intent $\phi_{t}^{\prime}$, we search $\mathcal{M}$ for the closest stored label:
@@
where $d(\cdot,\cdot)$ is the similarity metric used to determine whether the $\phi_{t}$ is semantically similar to any intent $\phi_{j}$. If $d\bigl{(}\phi_{t}^{\prime},\phi_{j^{*}}\bigr{)}\;\leq\;\epsilon_{\mathrm{ret}}$, we then retrieve and reuse the existing planner:
@@ IV-C Safety-critical Scenario Generation
Figure 5: Comparison of speed and acceleration distributions between raw (blue) and generated safety-critical (orange) trajectories.
@@
To validate the realism of these generated scenarios, we compared their kinematic distributions with the raw data, as shown in Fig. 5. The generated future trajectories remain statistically close to the original data, with low KL-Divergences for speed (0.141) and acceleration (0.138). The proportion of abnormal lateral acceleration ( $\>4m/s^{2}$) is only 0.17%. This demonstrates that while the generated scenarios exhibit more aggressive maneuvers to induce criticality, they adhere to physical constraints.
@@ V Conclusion
There are several directions for further enhancing interactive scenario generation. First, the uncertainty within the intent inference can be integrated. The current deterministic intent inference could be extended by leveraging the inherent stochasticity of LLMs. Second, extending the memory bank to support hierarchical intents and multi-agent coordination may yield more complex adversarial interactions. Third, while our current experiments are based on the Waymo dataset, the design of our framework allows for generalization, which should be extended to industry-standard simulators such as CARLA.